# A Survey on Computational Pathology Foundation Models: Datasets, Adaptation Strategies, and Evaluation Tasks

### v1

@@ Abstract
Computational pathology foundation models (CPathFMs) have emerged as a powerful approach for analyzing histopathological data, leveraging self-supervised learning to extract robust feature representations from unlabeled whole-slide images. These models, categorized into uni-modal and multi-modal frameworks, have demonstrated promise in automating complex pathology tasks such as segmentation, classification, and biomarker discovery. However, the development of CPathFMs presents significant challenges, such as limited data accessibility, high variability across datasets, the necessity for domain-specific adaptation, and the lack of standardized evaluation benchmarks. This survey provides a comprehensive review of CPathFMs in computational pathology, focusing on datasets, adaptation strategies, and evaluation tasks. We analyze key techniques, such as contrastive learning and multi-modal integration, and highlight existing gaps in current research. Finally, we explore future directions from four perspectives for advancing CPathFMs. This survey serves as a valuable resource for researchers, clinicians, and AI practitioners, guiding the advancement of CPathFMs toward robust and clinically applicable AI-driven pathology solutions.
@@ 1 Introduction
In recent years, foundation models (FMs) have gained significant attention in CPath Ochi et al. (2024). Unlike traditional deep learning models that require large-scale, manually labeled datasets for each specific task, computational pathology foundation models (CPathFMs) leverage self-supervised learning (SSL) to learn rich feature representations from unlabeled histopathological images. These models can be fine-tuned for various downstream pathology tasks using transfer learning, few-shot learning, or zero-shot learning, reducing the dependency on extensive expert manual annotations. Uni-modal CPathFMs are trained exclusively on histopathological images, capturing domain-specific features, while multi-modal CPathFMs, which integrate histopathological images with clinical linguistic data, further enhance AI-driven pathology by leveraging complementary information sources. Despite their success, effectively pre-training CPathFMs remains challenging due to data limitations, adaptation difficulties, and evaluation inconsistencies.
@@
While several survey papers have focused on CPathFMs, some of them emphasize benchmarking rather than a comprehensive investigation of existing methods Campanella et al. (2024a); Lee et al. (2024); Neidlinger et al. (2024). They cover too few approaches and lack a summary of pre-training datasets and evaluation tasks. Although Neidlinger et al., included a wide range of CPathFMs and pre-training datasets, their summarization of methods and datasets remains insufficiently detailed. Ochi et al., Ochi et al. (2024) and Chanda et al., Chanda et al. (2024) conducted surveys on a large number of CPathFMs, their pretraining datasets, and evaluation tasks. However, the methods mentioned by the former are relatively few and not up-to-date, while the latter did not provide a detailed introduction on how these methods are adapted to the pathology domain or distinguish between the different adaptations. Although both provided summaries of evaluation tasks, the latter only listed the tasks without offering a systematic taxonomy, whereas the former’s summary of the tasks was not comprehensive enough.
@@
  • Systematically reviewing adaptation techniques in pre-training CPathFMs, covering both uni-modal (image-based) and multi-modal (image and text) approaches, and highlighting their respective advantages and limitations.
@@
By synthesizing recent advancements and identifying open research challenges, this survey serves as a valuable resource for researchers, practitioners, and clinicians working at the intersection of CPath, medical AI, and foundation models.
@@ 2.1 Computational Pathology (CPath)
CPath is an interdisciplinary field that combines artificial intelligence, machine learning, and computer vision with digital pathology to enhance diagnosis, prognosis, and treatment planning. By leveraging whole-slide imaging and deep learning, CPath facilitates scalable and automated analysis of histopathological data, reducing reliance on manual examination by pathologists and improving diagnostic consistency.
Whole-slide imaging refers to the digital scanning of entire histopathology slides at high resolution, producing gigapixel WSIs that capture intricate tissue structures and morphologies. Due to their large size [^1] The size of WSIs varies depending on factors such as tissue type, magnification level, scanning resolution, and file format, but they are typically gigapixel-scale images and can range from hundreds of megabytes to several gigabytes per slide., WSIs are typically divided into smaller tile images, which serve as fundamental units for computational analysis. These tiles can be further annotated with tile captions, providing textual descriptions of specific tissue regions, while WSI reports summarize patient-level pathology interpretations (e.g., tumor type and grade).
Figure 1 presents an illustration of key histopathology data modalities, including WSIs and tile images at multiple magnifications (Field of View, FoV), and different imaging techniques, such as H&E, immunohistochemistry (IHC), and multiplex immunofluorescence (MxIF). These elements are essential for developing CPathFMs by incorporating multi-scale image representations and domain-specific knowledge. The FoV plays a critical role in image interpretation, with tile-level analysis focusing on localized cellular structures, while WSI-level analysis provides a broader view of tissue morphology. Multi-modal CPathFMs further integrate histopathological images with clinical reports and captions, improving model generalizability and interpretability in AI-assisted pathology.
@@
2.2 Self-Supervised Contrastive Learning
Self-supervised contrastive learning has become a key approach in pre-training CPathFMs, allowing models to learn rich representations from unlabeled histopathological images. These frameworks often (1) maximize the similarity between different augmented views of the same image and (2) minimize similarities between distinct images to train feature extractors. Several prominent methods have been adapted for CPathFM pre-training, mainly including DINO Caron et al. (2021), DINOv2 Oquab et al. (2023), CLIP Radford et al. (2021), and CoCa Yu et al. (2022).
The global modeling capability and scalability of ViT have driven recent research to focus on integrating ViT into SSL frameworks Caron et al. (2021). To address the instability during ViT training, DINO, which follows a student-teacher paradigm, performs self-distillation in a self-supervised manner. Specifically, it matches the output of the student network with the output of the teacher network after centering (an operation similar to batch normalization) and updates the teacher network using exponential moving average (EMA) He et al. (2020). DINOv2 enhances DINO by integrating iBOT Zhou et al. (2021), which employs Masked Image Modeling (MIM) as introduced by Masked Autoencoders (MAE) He et al. (2022). MIM works by randomly masking certain regions of an image and reconstructing the masked areas, enabling the model to learn valuable representations. DINOv2 effectively captures fine-grained cellular structures within broader tissue contexts, thereby improving generalization across diverse pathology datasets.
Beyond vision-only learning, multi-modal contrastive frameworks such as CLIP integrate image and text encoders, enabling zero-shot classification and cross-modal retrieval. CLIP enhances model interpretability for pathology AI by leveraging paired histopathological images and textual descriptions (e.g., pathology reports and expert annotations). CoCa extends CLIP by incorporating a multi-modal decoder that generates language signals from visual signals, combining contrastive learning with image captioning and report synthesis. This dual-tower architecture improves diagnostic reasoning in CPath by enabling both retrieval-based and generative tasks.
Table 1: Statistics of datasets used for pre-training CPathFMs.
@@ 2.2 Self-Supervised Contrastive Learning
Lu et al. (2024b) 1.18M Tile-Caption Pairs $\times×448$ - PMC OA EDU PathChat
467K Instructions - - PMC OA, TCGA Proprietary
@@ 2.3 Challenges in Pre-training CPathFMs
One of the primary challenges in pre-training CPathFMs is data availability. Histopathology datasets are often stored within institutional repositories or hospital databases, requiring ethical approvals and access permissions, which significantly limits the availability of large, diverse datasets. Publicly available datasets are scarce and frequently originate from a single institution, reducing data diversity. Additionally, WSIs are gigapixel-scale images, often reaching several gigabytes per file, which creates significant storage and computational burdens. Handling such large-scale data efficiently requires advanced data management, compression, and distributed storage solutions. Annotation is another major obstacle, as pathology image labeling requires expert knowledge, often involving multiple pathologists to ensure accuracy. This process is time-consuming and costly, leading to a shortage of well-annotated datasets. Furthermore, pathology data exhibit substantial variability in staining protocols, magnification levels, and organ-specific structures, introducing domain shifts that complicate model generalization. Data imbalance further exacerbates these issues, as rare disease types account for only a small fraction of available data, causing models to overfit common diseases while performing poorly on underrepresented conditions.
@@
Evaluation presents a significant challenge in pre-training CPathFMs. The wide range of downstream pathology tasks, including classification, retrieval, segmentation, survival prediction, and so on, complicates performance assessment, making it difficult to compare CPathFMs across different institutions and tasks. The heterogeneity in evaluation methodologies further hinders the establishment of a universal benchmarking framework, limiting the ability to systematically assess and compare model performance.
@@ 3 Pre-training Datasets in CPathFMs
Uni-modal CTransPath Wang et al. (2022) MoCo v3 Swin Transformer [28M] - Tiles S -
@@
GigaPath Xu et al. (2024) DINOv2 ViT-G/14 [1.1B] & LongNet [125M] - WSIs S, S -
@@
Virchow2 Zimmermann et al. (2024) DINOv2 ViT-B/16 [86M] - Tiles S -
@@
Multi-modal PLIP Huang et al. (2023) CLIP ViT-B/32 [86M] Transformer Layers [63M] Tiles D N
PathCLIP Sun et al. (2024) CLIP ViT-B/32 [86M] Transformer Layers [63M] Tiles D N
PathAsst ViT-B/32 [86M] Vicuna-13B [13B] Tiles F I
@@
PRISM Shaikovski et al. (2024) CoCa ViT-H/14 [632M] & Perceiver Net. [105M] BioGPT [345M, 172M] WSIs F, S N, F
CHIEF Wang et al. (2024) CLIP Swin Transformer [28M] Transformer Layers [63M] WSIs D N
PathChat Lu et al. (2024b) CoCa ViT-L/16 [307M] Llama 2 [13B] Tiles D I
@@
    Pre-training strategies: F: Frozen, S: From Scratch, D: Domain-Specific Tuning, I: Instruction Tuning, N: Non-instruction Tuning.
Although early CPathFMs used relatively small and homogeneous pre-training datasets, recent studies have shown that higher quality, larger scale, and more diverse pathology pre-training datasets are more beneficial for adapting the foundation models trained on natural image datasets or existing SSL frameworks to the pathology domain Zimmermann et al. (2024). Therefore, summarizing the datasets used for pre-training CPathFMs can provide valuable insights for future research on CPathFMs. Table 1 provides a summary of the key statistics of the pre-training datasets utilized by each method discussed in Section 4.
Most CPathFMs utilize multiple data sources to construct larger and more diverse pathology pre-training datasets. Aside from inaccessible private data, existing methods often acquire WSIs from large-scale public pathology datasets such as The Cancer Genome Atlas (TCGA) Weinstein et al. (2013), The Genotype-Tissue Expression (GTEx) Consortium Consortium et al. (2015), and the Pathology AI Platform (PAIP) Kim et al. (2021), as well as tile-caption pairs from the PubMed Central Open Access Dataset (PMC OA) Lu et al. (2024a). The diversity of data sources imposes higher requirements on data curation. Although the processes vary across methods, some steps are similar. As shown in Figure 2, images are typically subject to detection and segmentation of subfigures; resizing and cropping to meet model requirements, or using off-the-shelf LLMs to generate captions or reports for multimodal task needs. For text, titles are usually segmented or refined using LLMs. The cross-modal curation process includes the alignment of subfigures and subcaptions, as well as the construction of instructions. Finally, the data is filtered to select the required pathology data.
From the perspective of data types, the pre-training datasets for uni-modal CPathFMs are relatively straightforward, typically containing WSIs and the tiles extracted from them. However, multi-modal CPathFMs often use different data types as inputs depending on the training needs. Tile-caption pairs are typically used for tile-level, while WSI-report pairs are generally employed for slide-level (note that one or more WSIs usually have a corresponding clinical report in practice). However, Wang et al., Wang et al. (2024) uses the WSI labels (anatomical sites) as the textual information to construct the data pairs. For pathology LLM, instruction dataset are needed for fine-tuning, which consists of triplets that include a prompt specifying a given task, multi-modal supplementary information, and the corresponding desired output. Sun et al., Sun et al. (2024) developed two different types of instructions, one of which enables LLMs to invoke off-the-shelf pathology-specific sub-models to directly address pathology tasks, while Lu et al. , Lu et al. (2024b) utilized a diverse set of data sources from several institutions to create six types of instructions.
4 Adaptation Strategies in CPathFMs
SSL has been widely applied in the development of CPathFMs to address the lack of labels. These models typically adapt SSL frameworks that have proven successful in natural images and perform pre-training on carefully curated pathology datasets. Depending on the type of pathology data they used, these approaches can be categorized into uni-modal and multi-modal methods, which are summarized in Table 2.

### v2

@@ Abstract
Computational pathology foundation models (CPathFMs) have emerged as a powerful approach for analyzing histopathological data, leveraging self-supervised learning to extract robust feature representations from unlabeled whole-slide images. These models, categorized into uni-modal and multi-modal frameworks, have demonstrated promise in automating complex pathology tasks such as segmentation, classification, and biomarker discovery. However, the development of CPathFMs presents significant challenges, such as limited data accessibility, high variability across datasets, the necessity for domain-specific adaptation, and the lack of standardized evaluation benchmarks. This survey provides a comprehensive review of CPathFMs in computational pathology, focusing on datasets, adaptation strategies, and evaluation tasks. We analyze key techniques, such as contrastive learning, masked image modeling and multi-modal integration, and highlight existing gaps in current research. Finally, we explore future directions from four perspectives for advancing CPathFMs. This survey serves as a valuable resource for researchers, clinicians, and AI practitioners, guiding the advancement of CPathFMs toward robust and clinically applicable AI-driven pathology solutions.
@@ 1 Introduction
In recent years, foundation models (FMs) have gained significant attention in CPath Ochi et al. (2024). Unlike other deep learning models that require large amounts of labeled data and are limited to specific tasks, computational pathology foundation models (CPathFMs) use larger models (such as various ViTs) as the backbone and are pre-trained on a vast and diverse set of unlabeled histopathological data through self-supervised learning (SSL). These models can be fine-tuned for various downstream pathology tasks using transfer learning, few-shot learning, or zero-shot learning, reducing the dependency on extensive expert manual annotations. Uni-modal CPathFMs are trained exclusively on histopathological images, capturing domain-specific features, while multi-modal CPathFMs, which integrate histopathological images with clinical linguistic data, enhance AI-driven pathology by leveraging complementary information available in electronic health records (EHRs). Despite some striking early successes, effectively pre-training CPathFMs remains challenging due to data limitations, adaptation difficulties, and evaluation inconsistencies.
@@
While several survey papers have focused on CPathFMs, some of them emphasize benchmarking rather than a comprehensive investigation of existing methods Campanella et al. (2024a); Lee et al. (2024); Neidlinger et al. (2024). They cover too few approaches and lack a summary of pre-training datasets and evaluation tasks. Although Neidlinger et al., Neidlinger et al. (2024) included a wide range of CPathFMs and pre-training datasets, their summarization of methods and datasets remains insufficiently detailed. Ochi et al., Ochi et al. (2024) and Chanda et al., Chanda et al. (2024) conducted surveys on a large number of CPathFMs, but the methods they mentioned are not comprehensive and not up-to-date, and they did not provide a detailed introduction on how these methods are adapted to the pathology domain or distinguish between the different adaptations. Regarding summaries of evaluation tasks, the latter only listed the tasks without offering a systematic taxonomy, whereas the former’s summary of the tasks was not comprehensive enough.
@@
  • Systematically reviewing adaptation techniques in pre-training CPathFMs, covering 28 existing and up-to-date models across both uni-modal (image-based) and multi-modal (image-text) paradigms.
@@ 2.1 Computational Pathology (CPath)
CPath is an interdisciplinary field that combines artificial intelligence, machine learning, and computer vision with digital pathology to enhance diagnosis, prognosis, and treatment planning. By leveraging whole-slide imaging and deep learning, CPath facilitates scalable and automated analysis of histopathological data, reducing reliance on manual examination by pathologists and improving diagnostic consistency. However, despite dramatic progress in CPathFMs and CPath more broadly, the deployment of machine learning methods in real-world clinical settings remains challenging, with many limitations in performance, usability, and regulatory compliance yet to be overcome.
Whole-slide imaging refers to the digital scanning of entire histopathology slides at high resolution, producing gigapixel WSIs that capture intricate tissue structures and morphologies. Analysis of WSIs is standard practice during manual image review, and is required by the FDA (U.S. Food and Drug Administration) for digital pathology applications, Due to their large size [^1] The size of WSIs varies depending on factors such as tissue type and magnification level, scanning resolution, and file format, but they are typically gigapixel-scale images and can range from hundreds of megabytes to several gigabytes per slide., computational analysis WSIs typically involves dividing them into smaller tile images. WSI reports summarize patient-level pathology interpretations and patient level data (e.g., tumor type and grade) but individuals tiles can also be annotated with tile captions, providing textual descriptions of specific tissue regions.
Figure 1 presents an illustration of key histopathology data modalities, including WSIs and tile images at multiple magnifications (Field of View, FoV), and different imaging techniques, such as H&E, immunohistochemistry (IHC), and multiplex immunofluorescence (MxIF). H&E and IHC images are routinely used in clinical histopathology services and diagnosis but MxIF is a research technique still in development for diagnostic and clinical applications. All three types of images have been used in CPathFMs. The FoV plays a critical role in image interpretation, with tile-level analysis focusing on localized cellular structures, while WSI-level analysis provides a broader view of tissue morphology. Multi-modal CPathFMs further integrate histopathological images with clinical reports and captions, improving model generalizability and interpretability in AI-assisted pathology.
@@
2.2 ViT-Based SSL Frameworks for FMs
The ViT has emerged as a cornerstone for foundation models due to its powerful self-attention mechanism, which excels in capturing global dependencies. Its patch-based tokenization is inherently compatible with masked learning strategies, making it a natural fit for SSL. Furthermore, ViT’s scalability aligns with the large-scale nature of foundation models, and its Transformer architecture seamlessly integrates multiple modalities, such as vision and language. These properties have led to the development of various SSL frameworks that effectively leverage ViT, laying the groundwork for foundation models in pathology and beyond.
A key approach in SSL for ViTs is Masked Image Modeling (MIM), which is introduced by Masked Autoencoders (MAE) He et al. (2022). MIM leverages the patch-based structure of ViT to predict missing patches, thereby effectively learning rich and robust representations. Inspired by the success of BERT in NLP, BEiT Bao et al. (2021) incorporates a discrete tokenizer in MIM, while BEiT v2 Peng et al. (2022) refines the tokenization process through knowledge distillation based on BEiT.
In parallel, contrastive learning has played a crucial role in ViT-based SSL. DINO Caron et al. (2021), which follows a student-teacher paradigm, introduces a self-distillation approach that leverages ViT’s global attention to learn robust features without explicit labels. DINOv2 Oquab et al. (2023) enhances DINO by incorporating iBOT Zhou et al. (2021), which combines MIM with self-distillation. This integration allows DINOv2 to effectively capture fine-grained structures, improving its generalization across diverse datasets, and positioning it as a strong candidate for FMs.
Beyond vision-only learning, multi-modal contrastive frameworks such as CLIP Radford et al. (2021) integrate image and text encoders, enabling zero-shot classification and cross-modal retrieval. CoCa Yu et al. (2022) extends CLIP by incorporating a multi-modal decoder that generates language signals from visual signals, combining contrastive learning with image captioning and report synthesis. In terms of MIM, BEiT-3 Wang et al. (2022a) builds upon BEiT v2 by introducing more advanced token prediction mechanisms and leveraging multi-modal learning to extend its capabilities
Table 1: Statistics of pathology datasets used for pre-training CPathFMs. The “-” symbol represents the absence of relevant information.
@@ 2.2 ViT-Based SSL Frameworks for FMs
Ma et al. (2024) 72.3K 190.2M $\times×512$ - TCGA, GTEx, etc. - GPFM
[…]
Saillard et al. (2024) 500K+ 100M+ - - - Proprietary H-optimus-0
@@
Zhou et al. (2024b) A KG with 50.5K Pathology Attributes - - - - - OncoTree, etc. - KEP
576.6K, 138.9K Tile-Caption Pairs $\times×224$ - - Quilt-1M, OpenPath -
@@
Zhou et al. (2024a) A KG with 139K Disease Attributes - - - - - DO, UMLS - KEEP
143K Tile-Caption Pairs $\times×224$ - - Quilt-1M, OpenPath -
[…]
@@ 2.3 Challenges in Pre-training CPathFMs
A primary challenge in pre-training CPathFMs is data availability. Histopathology datasets are often restricted by ethical approvals and access permissions, limiting the availability of large, diverse datasets. Public datasets are rare, often from a single institution, reducing diversity. Additionally, WSIs are large, gigapixel-scale images that pose storage and computational challenges, requiring advanced data management solutions. Annotation is another major obstacle, as pathology image labeling requires expert knowledge, often involving multiple pathologists to ensure accuracy. This process is time-consuming and costly, leading to a shortage of well-annotated datasets. Furthermore, pathology data exhibit substantial variability in staining protocols, magnification levels, and organ-specific structures, introducing domain shifts that complicate model generalization. Data imbalance further exacerbates these issues, as rare disease types account for only a small fraction of available data, causing models to overfit common diseases while performing poorly on underrepresented conditions.
@@
Evaluation presents a significant challenge in pre-training CPathFMs. The wide range of downstream pathology tasks, including classification, retrieval, segmentation, generation, and so on, complicates performance assessment, making it difficult to compare CPathFMs across different institutions and tasks. The heterogeneity in evaluation methodologies further hinders the establishment of a universal benchmarking framework, limiting the ability to systematically assess and compare model performance.
@@ 3 Pre-training Datasets in CPathFMs
Although early CPathFMs used relatively small and homogeneous pre-training datasets, recent studies have shown that higher quality, larger scale, and more diverse pathology pre-training datasets are more beneficial for adapting the foundation models trained on natural image datasets or existing SSL frameworks to the pathology domain  Zimmermann et al. (2024). Therefore, summarizing the datasets used for pre-training CPathFMs can provide valuable insight into requirements for future research on CPathFMs. Table 1 provides a summary of pre-training datasets utilized by each method discussed in Section 4.
Most CPathFMs utilize multiple data sources to construct larger and more diverse pathology pre-training datasets. Aside from inaccessible private data, existing methods often acquire WSIs from large-scale public pathology datasets such as The Cancer Genome Atlas (TCGA) Weinstein et al. (2013), The Genotype-Tissue Expression (GTEx) Consortium Consortium et al. (2015), as well as tile-caption pairs from the PubMed Central Open Access Dataset (PMC OA) Lu et al. (2024). A common method for multi-modal pathology data is to collect it from the internet, for example, Ikezogwo et al., Ikezogwo et al. (2024) curate a large-scale dataset named Quilt-1M from multiple different online sources. Huang et al., Huang et al. (2023) and Sun et al., Sun et al. (2024) also released their pretraining datasets, OpenPath and PathCap, respectively, alongside the proposal of CPathFM.
The diversity of data sources imposes higher requirements on data curation. Although the processes vary across methods, some steps are similar. As shown in Figure 2, images are typically subject to detection and segmentation of subfigures; resizing and cropping to meet model requirements, or using off-the-shelf LLMs to generate captions or reports for multimodal task needs. For text, titles are usually segmented or refined using LLMs. The cross-modal curation process includes the alignment of subfigures and subcaptions, as well as the construction of instructions. Finally, the data is filtered to select the required pathology data.
From the perspective of data types, the pre-training datasets for uni-modal CPathFMs are relatively straightforward, typically containing WSIs and the tiles extracted from them. However, multi-modal CPathFMs often use different data types as inputs depending on the training needs. Tile-caption pairs are typically used for tile-level, while WSI-report pairs are generally employed for slide-level (note that one or more WSIs usually have a corresponding clinical report in practice). Notably, Wang et al., Wang et al. (2024) used the WSI labels (anatomical sites) as the textual information to construct the data pairs.
4 Adaptation Strategies in CPathFMs
@@
Uni-modal CTransPath Wang et al. (2022b) MoCo v3* Swin Transformer [28M] - Tiles S -
@@
PLUTO Juyal et al. (2024) DINOv2* FlexiViT-S [22M] - Tiles S -
GigaPath Xu et al. (2024) DINOv2* ViT-G/14 [1.1B] & LongNet [125M] - WSIs S, S -
@@
BEPH Yang et al. (2024) BEiTv2 ViT-B/16 [86M] & VQ-KD [86M] - Tiles D, D -
GPFM Ma et al. (2024) DINOv2* ViT-L [307M] - Tiles S -
Virchow2 Zimmermann et al. (2024) DINOv2* ViT-B/16 [86M] - Tiles S -
@@
H-optimus-0 Saillard et al. (2024) DINOv2 ViT-G/14 [1.1B] - Tiles S -
Atlas Alber et al. (2025) DINOv2 ViT-H/14 [632M] - Tiles D -
Multi-modal PLIP Huang et al. (2023) CLIP ViT-B/32 [86M] Transformer Layers [63M] Tiles D D
PathCLIP Sun et al. (2024) CLIP ViT-B/32 [86M] Transformer Layers [63M] Tiles D D
QuiltNet Ikezogwo et al. (2024) CLIP ViT-B/(16,32) [86M] GPT-2 [1.5B] & PubMedBERT [100M] Tiles D D
@@
PRISM Shaikovski et al. (2024) CoCa ViT-H/14 [632M] & Perceiver Net. [105M] BioGPT [345M, 172M] WSIs F, S D, F
CHIEF Wang et al. (2024) CLIP* Swin Transformer [28M] Transformer Layers [63M] WSIs D D
KEP Zhou et al. (2024b) CLIP* ViT-B/(16,32) [86M] PubMedBERT [100M] Tiles D S
@@
KEEP Zhou et al. (2024a) CLIP* ViT-L [307M] PubMedBERT [100M] Tiles D S
MUSK Xiang et al. (2025) CoCa*, BEiT-3 V-FFN [202M] L-FFN [202M] Tiles S S
Shared Attention Layers [202M]
    Made domain-specific improvements or extensions to the SSL framework for pathology.
@@
    Pre-training strategies: F: Frozen, S: From Scratch, D: Domain-Specific Tuning.
SSL has been widely applied in the development of CPathFMs to address the lack of labels. These models typically adapt SSL frameworks that have proven successful in natural images, and perform pre-training on carefully curated pathology datasets. Depending on the type of pathology data they used, these approaches can be categorized into uni-modal and multi-modal methods, as introduced in Table 2.